OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (lgbasallote):

volume of a cone = \(\frac{1}{3} \pi r^2 h\) where do you have troubles?

OpenStudy (anonymous):

i cant find the height

OpenStudy (lgbasallote):

okay...start by identifying the values given to you...which of those are given?

OpenStudy (anonymous):

the overall volume wich is 300 cubic km and the radius wich is 9 km

OpenStudy (lgbasallote):

good! and it says that pi is 3.14 so we substitute... \[300 = \frac{1}{3} (3.14)(9)^2h\] good?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

1/3 x 3.14 x 81 x h = 300 cubic km

OpenStudy (anonymous):

now how should i find the height

OpenStudy (lgbasallote):

good \[300 = \frac{1}{3} (3.14) (81) h\] first we multiply both sides by 3 to get rid of 1/3

OpenStudy (anonymous):

ok so that would be 9.42 x 243?

OpenStudy (lgbasallote):

no no no...if we multiply the right side by 3...then 1/3 just cancels out okay?

OpenStudy (lgbasallote):

\[\frac{1}{3} \times 3 = 1\]

OpenStudy (anonymous):

i did not get that...

OpenStudy (lgbasallote):

\[\frac{1}{3} (3.14)(81) h \times 3 = \frac{3}{3} (3.14)(81) h\] how about that?

OpenStudy (anonymous):

ok so i multiply 1/3 by 3 not anyting else?

OpenStudy (lgbasallote):

nothing else on the right side..you multiply 3 to 300 as well

OpenStudy (anonymous):

i still dont understand well im sorry

OpenStudy (lgbasallote):

according to multiplicative property of equality..you need to multiply BOTH sides by THE SAME number \[300 = \frac{1}{3} (3.14)(81) h \rightarrow 300(3) = \frac{3}{3} (3.14)(81)h\] okay?

OpenStudy (lgbasallote):

i just multiply BOTH sides by 3

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

so what's 300 x 3?

OpenStudy (anonymous):

900

OpenStudy (lgbasallote):

good and what's \(\large \frac{3}{3}\)

OpenStudy (anonymous):

it is 3

OpenStudy (lgbasallote):

it's not 3...that's like 3 divided by 3

OpenStudy (lgbasallote):

a number divided by itself...

OpenStudy (anonymous):

then its 0 ?

OpenStudy (anonymous):

no sorry its 1

OpenStudy (lgbasallote):

final answer /:)

OpenStudy (anonymous):

so my height is 1?

OpenStudy (lgbasallote):

no...im asking if you're sure that \(\large \frac{3}{3} = 1\)

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then it becomes \[900 = (3.14)(81) h\] agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what's \(3.14 \times 81?\) Hint: use a calculator

OpenStudy (anonymous):

254.34

OpenStudy (lgbasallote):

good..now we divide both sides by 254.34 \[\large \frac{900}{254.34} = \frac{(\cancel{254.34})h}{\cancel{254.34}}\]

OpenStudy (anonymous):

ok

Parth (parthkohli):

A 9 kilometer radius of the base. A huge cone. Seriously, people who make textbooks don't ever think for real.

OpenStudy (anonymous):

3.53857042

OpenStudy (lgbasallote):

there we go!

Parth (parthkohli):

Round to the nearest tenths @mumina1999

OpenStudy (lgbasallote):

dont forget units ;D

OpenStudy (anonymous):

ok thanks guys